# Proton decay in de Sitter environment

### v3

@@ II.1 Ionization of hydrogen atom
$r<r_{0}=\frac{1}{H}\sqrt{\frac{2\epsilon}{m_{e}}}\ll r_{H}=\frac{1}{H}\,,$ (7)
@@
In Eq.(4) we ignore the pre-exponential factor. In this approximation the decay rate coincides with that, which comes from the activation process in the Minkowski state in the presence of the heat bath with temperature $T=H/\pi$. This demonstrates that matter perceives the de Sitter environment as the heat bath with temperature $T=H/\pi$. The local temperature of the de Sitter heat bath governs also the other processes of the splitting of stable objects, including the decay of proton in de Sitter environment.
@@ III.1 Decay of proton in de Sitter environment
III.2 Proton as the source of electron-positron pairs
@@ IV Vacuum back reaction and de Sitter decay
On the other hand, the de Sitter vacuum can be considered as the combination of dark energy with the equation of state $w=-1$ and dark matter with the equation of state $w=+1$, which is the analog of the Zel’dovich stiff matter.Volovik2024ad

### v4

@@ II.1 Ionization of hydrogen atom
$r<r_{0}=\frac{1}{H}\sqrt{\frac{2\epsilon}{m_{e}}}\ll r_{H}=\frac{1}{H}\,.$ (7)
Then one has
[…]
@@
In Eq.(4) we ignore the pre-exponential factor. In this approximation the decay rate coincides with that, which comes from the activation process in the Minkowski state in the presence of the heat bath with temperature $T=H/\pi$. This demonstrates that matter perceives the de Sitter environment as the heat bath with temperature $T=H/\pi$. The local temperature of the de Sitter heat bath governs also the other processes of the splitting of stable objects, including the decay of proton in de Sitter environment. It is important that due to the de Sitter symmetry, the local temperature is the same for all points in the de Sitter space, i.e. for all observers, that are static in the de Sitter environment. The coordinate transformation in the PG metric (1) which connects two static observers:
${\bf r}\rightarrow{\bf r}+{\bf a}\,e^{Ht}\,\,,\,\,t\rightarrow t\,.$ (9)
The metric (1) is invariant under this shift of the position of the observer. Each observer considers the hydrogen atom, which is at rest in his/her reference frame.
@@ III.1 Decay of proton in de Sitter environment
III.2 Decay of proton in the Schwarzschild-de Sitter environment
We can use the conventional static metric for the Schwarzschild-de Sitter (SdS) black hole:
$ds^{2}=-\left(1-\frac{2M}{r}-H^{2}r^{2}\right)dt^{2}+\frac{dr^{2}}{1-\frac{2M}{r}-H^{2}r^{2}}+r^{2}d\Omega^{2}\,.$ (22)
Let us place the proton at the point $r=r_{0}$, where
$r_{0}^{3}=\frac{GM}{H^{2}}\,.$ (23)
The spherical surface at $r=r_{0}$ is the zero-gravity surface, and thus we can find the proton decay rate, when proton is rest. Neat this point, at $|r-r_{0}|\ll r_{0}$, the metric is
$ds^{2}=-\left(1-3H^{2}(r-r_{0})^{2}\right)dt^{2}+\frac{dr^{2}}{1-3H^{2}(r-r_{0})^{2}}+r^{2}d\Omega^{2}\,.$ (24)
The decay rate in such environment corresponds to the decay rate in the de Sitter state with the effective Hubble parameter $\tilde{H}=\sqrt{3}H$. Then the temperature, which determines the decay of proton in the SdS environment, is (see also Ref. Volovik2023 ):
$T_{SdS}=\frac{\sqrt{3}}{\pi}H\,.$ (25)
It is interesting that this temperature does not depend on the mass of black hole and is fully determined by the Hubble parameter. The process of decay takes place in the region between the black hole horizon and the cosmological horizon, $r_{-}<r_{0}<r_{+}$, and thus it is the local process, which has no relation to the Hawking radiation from the horizons. However, in the limit when the two horizons merge we can compare this activation temperature with the Bousso-Hawking temperature.BoussoHawking1996 In this Nariai limit the Bousso-Hawking temperature $T_{\rm BH}$ is twice smaller than the temperature which determines the proton decay in the SdS spacetime, $T_{\rm BH}=\frac{\sqrt{3}}{2\pi}H=T_{SdS}/2$.
The quantum tunneling approach also determines the thermodynamics of black holes. The quantum tunneling process of radiation of particles determines the Hawking temperature, while the quantum tunneling of macroscopic objects (the macroscopic quantum tunneling) determines the non-extensive entropy of the black hole. Volovik2024TC
III.3 Proton as the source of electron-positron pairs
@@ IV Vacuum back reaction and de Sitter decay
On the other hand, the pure de Sitter vacuum can be considered as the combination of dark energy with the equation of state $w=-1$ and dark matter with the equation of state $w=+1$.Volovik2024ad The dark matter component of the de Sitter vacuum has the same equation of state as the Zel’dovich stiff matter.Zeldovich1962